# DR-MPC: Deep Residual Model Predictive Control for Real-world Social Navigation

### v1

@@ Abstract
How can a robot safely navigate around people exhibiting complex motion patterns? Reinforcement Learning (RL) or Deep RL (DRL) in simulation holds some promise, although much prior work relies on simulators that fail to precisely capture the nuances of real human motion. To address this gap, we propose Deep Residual Model Predictive Control (DR-MPC), a method to enable robots to quickly and safely perform DRL from real-world crowd navigation data. By blending MPC with model-free DRL, DR-MPC overcomes the traditional DRL challenges of large data requirements and unsafe initial behavior. DR-MPC is initialized with MPC-based path tracking, and gradually learns to interact more effectively with humans. To further accelerate learning, a safety component estimates when the robot encounters out-of-distribution states and guides it away from likely collisions. In simulation, we show that DR-MPC substantially outperforms prior work, including traditional DRL and residual DRL models. Real-world experiments show our approach successfully enables a robot to navigate a variety of crowded situations with few errors using less than 4 hours of training data (video: https://youtu.be/GUZlGBk60uY, code: upon acceptance).
@@
Social HRI, Reinforcement Learning, Model Predictive Control, Real-world Robotics, Motion Planning.
@@ I Introduction
Achieving reliable robotic navigation remains one of the main challenges in integrating mobile robots into everyday environments. Applications such as food service, warehouse logistics, and equipment transport could see substantial cost and time savings but require robots to navigate dense human environments with static obstacles.
@@
DRL-based social navigation offers a compelling value proposition: learning efficient and safe robot behaviour without relying on an accurate human behaviour model. Most DRL social navigation research centers around simulation due to the dangers posed by randomly initialized DRL agents and the extensive data required for training [5]. However, these simulators contain human models that are inevitably mismatched with the true dynamics of human motion. Thus, agents trained in these simulators are unfit for deployment. For instance, the popular CrowdNav simulator assumes a cooperative and deterministic human policy [6]: Optimal Reciprocal Collision Avoidance (ORCA) [7]. This assumption leads DRL agents to learn overly aggressive behaviours [8]. Additionally, many simulators such as [6] neglect the presence of static obstacles, further exacerbating the sim-to-real gap.
@@
In this paper, we train a DRL-based social navigation model directly in the real world. To handle environments with static obstacles, we modify the typical DRL social navigation Markov Decision Process (MDP) into human avoidance with path tracking within virtual corridors. We introduce a novel approach, Deep Residual Model Predictive Control (DR-MPC), that integrates Model Predictive Control (MPC) path tracking to significantly accelerate the learning process. Lastly, we design a pipeline with out-of-distribution (OOD) state detection and a heuristic policy to guide the DRL agent into higher-reward regions. Our approach enabled real-world deployment of a DRL agent without any simulation with less than 4 hours of data.
We begin by discussing related work. Then, we define our decision process formulation in Section III. Next, in Section IV, we detail and motivate our design of DR-MPC. Then, in Section V, we explain how DR-MPC fits into the overall pipeline. In Section VI, we quantitatively demonstrate the effectiveness of our approach in simulation compared to other models. Finally, in Section VII, we present an analysis of our algorithm’s performance when trained on a real robot.
@@ II Related Works
II-A Social Navigation Simulators
Several 2D and 3D simulators have been developed and continue to be developed for social navigation. These simulators offer a way to safely test and experiment with different approaches. Simulators also offer the possibility, in theory, of developing ‘sim-to-real’ approaches, which can substantially reduce the amount of real-world learning (and mistakes) that robots make. A popular simulator for social navigation is the 2D CrowdNav simulator [6], designed for waypoint navigation in open spaces with the human policy defined by ORCA [7]. While CrowdNav has facilitated substantial DRL model development [9, 8, 10, 11], the cooperative nature of the ORCA policy creates a large sim-to-real gap. When the robot is visible to humans, the learned DRL policy is dangerously aggressive: the robot pushes humans out of its way to reach its goal [8]. Consequently, the invisible testbed—when the robot is invisible to the humans—was adopted as the benchmark standard. Although the invisible testbed reduces the DRL agent’s aggression, the problem becomes equivalent to decoupled planning where the DRL agent does not learn how its action will influence humans [8].
Looking beyond CrowdNav, other simulators use alternative human motion models, such as variations of the Social Forces Model (SFM) or behaviour graphs [12, 13, 14]. Unfortunately, even the latest simulators struggle with human realism and often exhibit unsmooth human motions and enter deadlock scenarios. These deficiencies create a significant sim-to-real gap, where models trained in simulation often perform differently–—and usually worse–—when applied in the real world [15]. While these simulators have been essential tools for advancing DRL model architectures, the most accurate data for DRL social navigation is real-world data. By reducing the amount of training data required, we enable the direct training of DR-MPC in the real world, avoiding unnecessary inductive biases introduced by simulators.
@@ II-B RL Social Navigation Models
Crowd navigation, a major sub-field of DRL social navigation, focuses on enabling a robot to navigate among humans in open areas. Significant progress has been made in incorporating machine learning advancements to enhance model architectures that enable reasoning about crowd dynamics in both spatial and temporal dimensions.
@@
Early models analyzed the scene by considering humans individually [16, 17]. Then, approaches incorporated attention mechanisms to reason about the crowd as a whole, using learned attention scores to generate a crowd embedding for decision making [6, 18, 19]. Subsequent advances analyzed both human-robot and human-human interactions through graph neural networks (GNNs) [9, 19, 10]. Most recently, models include temporal reasoning using spatio-temporal graphs, multihead attention mechanisms, and transformers [8, 11], enabling reasoning on the trajectory level.
@@
This paper does not focus on developing a novel architecture for traditional waypoint-based crowd navigation. Instead, we leverage state-of-the-art (SOTA) architectures for processing human-related inputs and incorporate them into DR-MPC.
II-C Residual DRL
@@
Residual DRL integrates two policies together: one is user-supplied, another is learned [20]. DRL learns corrective actions on top of this base controller. The learning speedup occurs because on initialization, the DRL agent will have an expected action equal to the base controller, which, assuming a suitable user-supplied controller, can result in much better performance than a random policy [20].
@@
To the best of our knowledge, residual DRL has not been applied to social navigation, but some related works have attempted to incorporate classical control. For example, Kästner et al. [21] learn a DRL policy that switches between different controllers, each optimized for an individual task. Another work, Semnani et al. [22] switch from a DRL policy to a force-based model when close to humans. While these approaches reduce the learning burden on DRL, they limit the overall performance by the capabilities of the base controllers. When the overall policy selects a base controller, DRL can not optimize the performance of these individual controllers.
@@
In contrast, DR-MPC fully exploits the capabilities of MPC path tracking. DR-MPC can either replicate the MPC action or generate an action that diverges significantly from it. Unlike residual DRL, which follows the base controller actions in expectation, DR-MPC’s initial behaviour almost exactly follows MPC path tracking, representing the best achievable performance without any prior information of humans. Additionally, DR-MPC dynamically learns when to integrate or disregard the MPC action, leading to significantly faster training compared to traditional residual DRL.
II-D Path Tracking
@@
DRL is beneficial for path tracking when modeling difficulties arise, whether related to the agent or environmental disturbances. However, in indoor environments with minimal disturbances and a typical robot that can be accurately modeled with unicycle kinematics, we can leverage MPC for path tracking using the formulation presented by Sehn et al. [23]. As a result, DRL does not need to explicitly learn path tracking, thereby reducing the overall learning complexity.
@@ III Decision process formulation
III-1 State Space
We construct our state $\mathcal{S}=\{\mathcal{S}_{\text{PT}},\mathcal{S}_{\text{HA}}\}$, where $\mathcal{S}_{\text{PT}}$ is the state information for path tracking and $\mathcal{S}_{\text{HA}}$ is the state information for human avoidance.
As in [8], we exclude human velocities due to the difficulty of estimating these quantities in the real world and assume a constant human radius. If at time $t$ there are $n_{t}$ visible humans: $\mathcal{S}_{\text{HA}}=\left\{\mathbf{v}^{t-H:t-1},\mathbf{r}^{t-H:t},\mathbf{q}^{t-H_{1}:t}_{1},\dots,\mathbf{q}^{t-H_{n_{t}}:t}_{n_{t}}\right\}$, where $\mathbf{v}^{t-H:t-1}$ is the robot’s past velocities from time $t-H$ to $t-1$, $\mathbf{r}^{t-H:t}$ is the robot’s past positions in the current robot frame from time $t-H$ to $t$, and $\mathbf{q}^{t-H_{i}:t}_{i}=\begin{bmatrix}\mathbf{p}^{t-H_{i}}_{i}\dots\mathbf{p}^{t}_{i}\end{bmatrix}$ is the positional trajectory of the $i^{\text{th}}$ visible human in the current robot frame with a look-back of $H_{i}$, capped at length $H$.
We select a local path representation as in [24] to generalize to any path length. Given that $\mathbf{p}_{\ell}$ is the closest localized node to the robot, $\mathcal{S}_{\text{PT}}=[\mathbf{p}_{\ell-L}\ldots\mathbf{p}_{\ell}\ldots\mathbf{p}_{\ell+F}]$, where $L$ is the look-back and $F$ is the look-forward parameter. These path nodes are also transformed into the current robot frame.
III-2 Actions
Our DRL agent interacts at the level of velocity. Our action space is a linear and angular velocity: $\mathbf{a}=\begin{bmatrix}v\hskip 5.69046pt\omega\end{bmatrix}$.
III-3 Rewards
Our reward function considers path advancement, path deviation, goal reaching, corridor collisions, small speeds, human collisions, and human disturbance:
@@ III-3 Rewards
where rewards marked with an asterisk are terminal rewards; on all termination conditions, the robot stops moving. All the following $\kappa$ are tuning parameters where $\kappa>0$. Path advancement: $r_{\text{pa}}=\kappa_{\text{pa}}(s^{\prime}_{i}-s_{i})$ where $s_{i}$ and $s^{\prime}_{i}$ are the arclength positions along the path at $\mathcal{S}$ and $\mathcal{S}^{\prime}$. Deviation: $r_{\text{dev}}=\kappa^{\text{xy}}_{\text{dev}}d_{\text{xy}}+\kappa^{\theta}_{\text{dev}}|d_{\theta}|$ where $d_{\text{xy}}$ is the Euclidean distance and $d_{\theta}$ is the angular offset from the closest point on the path. Goal: $r_{\text{goal}}^{*}=-\kappa_{\text{goal}}$ if the heading difference exceeds a threshold and 0 otherwise. Corridor collision: $r_{\text{cor-col}}^{*}=-\kappa_{\text{cor-col}}$ for colliding with the corridor. Minimal actuation: $r_{\text{act}}^{*}=-\kappa_{\text{act}}$ if the sum of the robot’s past $H$ speeds falls below a threshold.
For human avoidance, we align our rewards with the two most important principles of human-robot interaction (HRI): safety and comfort [25]. For safety, human collision: $r_{\text{hum-col}}^{*}=-\kappa_{\text{hum-col}}.$ For comfort, disturbance penalty: $r_{\text{dist}}=-\sum_{i=1}^{n_{t}}\left(\kappa_{\text{v}}|\Delta v_{\text{h}}^{i}|+\kappa_{\theta}|\Delta\theta_{\text{h}}^{i}|\right)$, which penalizes the robot for causing changes in a human’s velocity ($\Delta v_{\text{h}}^{i}$) and heading direction ($\Delta\theta_{\text{h}}^{i}$). These changes are computed relative to the human’s velocity and heading at the previous time step.
@@ IV DR-MPC Policy Architecture
Our DR-MPC policy architecture (Figure 3) consists of two main components: individual processing of $\mathcal{S}_{\text{PT}}$ and $\mathcal{S}_{\text{HA}}$, and the information fusion to generate a single action.
For processing $\mathcal{S}_{\text{PT}}$, we employ a Multi-layer Perceptron (MLP) to generate a latent embedding of the path information. Also, an MPC optimization guided by [23] generates $\mathbf{a}_{\text{MPC}}$.
To process $\mathcal{S}_{\text{HA}}$, we modify [10] to handle varying-length human trajectories for off-policy learning. We refer to this adapted architecture as the Human Avoidance Network (HAN); further details are provided in the appendix. HAN is a modular component of our pipeline and can be replaced with other crowd navigation work.
The output of HAN is a ‘crowd embedding’ that is used to generate six candidate human avoidance actions ($\mathbf{a}_{\text{HA}}$), where each $\mathbf{a}^{\text{i}}_{\text{HA}}$ is positioned in a different cell of the action space. The human-avoidance action space is partitioned into a grid of six cells defined by two linear velocity bins $[v_{\text{min}},v_{\text{middle}}]$, $[v_{\text{middle}},v_{\text{max}}]$, and three angular velocity bins $[w_{\text{min}},w_{\text{lower}}]$, $[\omega_{\text{lower}},\omega_{\text{upper}}]$, $[\omega_{\text{upper}},\omega_{\text{max}}]$. The inclusion of multiple actions is motivated by the fact that, in many cases, several viable actions may exist for human avoidance. We found empirically that this design reduces learning time compared to having the model learn the diversity. The output of the MLP following HAN is the mean action within each cell. Using a predetermined variance that decays over time, we sample a Gaussian to get $\mathbf{a}^{\text{i}}_{\text{HA}}$; this standard formulation comes from [26].
@@
Unlike residual DRL, which follows the base controller in expectation, DR-MPC begins with near-MPC behaviour by biasing actions toward MPC, initializing $\beta=-0.8$ to suppress $\boldsymbol{\alpha}$. Note that initializing $\beta$ too low yields little qualitative difference and requires a lot of model updates to raise $\beta$ towards 0. We observe the model naturally learns to adjust $\beta$ towards 0 to be able to take non-MPC actions.
Although DR-MPC’s effective action space after fusion is smaller than regular DRL, it contains the best actions for path tracking and human avoidance. While DRL with the full action space might theoretically achieve higher performance given infinite training, it faces practical challenges in balancing exploration and exploitation. Despite having a reduced action space, we can still achieve a high-performing policy but with significantly faster learning times.
The DRL algorithm we employ to train our model is Discrete Soft Actor-Critic (DSAC) with double average clipped Q-learning with Q-clip to properly backpropagate through $\mathbf{p}$ [28]. We further augment DSAC with the entropy formulation from SAC to optimize the log-standard deviation that generates $\alpha^{i}_{\text{raw}}$ [27]. Also note that Figure 3 represents the actor model within our actor-critic framework. The critic model shares the same architecture up to and including the first MLP in the fusion stage; however, this MLP outputs the Q-value. The action is included as an input to this MLP for the critic.

### v2

@@ Abstract
How can a robot safely navigate around people with complex motion patterns? Deep Reinforcement Learning (DRL) in simulation holds some promise, but much prior work relies on simulators that fail to capture the nuances of real human motion. Thus, we propose Deep Residual Model Predictive Control (DR-MPC) to enable robots to quickly and safely perform DRL from real-world crowd navigation data. By blending MPC with model-free DRL, DR-MPC overcomes the DRL challenges of large data requirements and unsafe initial behavior. DR-MPC is initialized with MPC-based path tracking, and gradually learns to interact more effectively with humans. To further accelerate learning, a safety component estimates out-of-distribution states to guide the robot away from likely collisions. In simulation, we show that DR-MPC substantially outperforms prior work, including traditional DRL and residual DRL models. Hardware experiments show our approach successfully enables a robot to navigate a variety of crowded situations with few errors using less than 4 hours of training data (video: https://youtu.be/GUZlGBk60uY, code: https://github.com/James-R-Han/DR-MPC).
@@
Social HRI, Reinforcement Learning, Model Predictive Control, Real-world Robotics, Autonomous Agents.
@@ I Introduction
Achieving reliable robotic navigation remains one of the main challenges in integrating mobile robots into society. Applications such as food service and equipment transport could see major cost and time savings but require robots to navigate human environments with static obstacles.
@@
DRL offers a compelling value proposition: learn efficient and safe robot behaviour without a human motion model. Current research centers around simulation due to the dangers posed by randomly initialized DRL agents and the extensive training data requirement [5]. However, these simulators use human models that are mismatched with the true human motion, making sim-to-real agents unfit for deployment. For instance, the popular CrowdNav simulator assumes a cooperative and deterministic human policy, [6] Optimal Reciprocal Collision Avoidance (ORCA) [7], which leads to aggressive DRL agents [8]. Also, many simulators neglect the presence of static obstacles, further exacerbating the sim-to-real gap.
@@
In this paper, we train a DRL agent directly in the real world. To handle environments with static obstacles, we modify the typical DRL social navigation Markov Decision Process (MDP) into human avoidance with path tracking within virtual corridors. We introduce a novel approach, Deep Residual Model Predictive Control (DR-MPC), that integrates Model Predictive Control (MPC) path tracking to significantly accelerate the learning process. Lastly, we design a pipeline with out-of-distribution (OOD) state detection and a heuristic policy to guide the DRL agent into higher-reward regions. Our approach enabled real-world deployment of a DRL agent without any simulation with less than 4 hours of data.
@@ II Related Works
Social Navigation Simulators. Several 2D and 3D simulators continue to be developed for social navigation. These simulators offer a way to safely test and experiment with different approaches. Simulators also offer the possibility, in theory, of developing ‘sim-to-real’ approaches, which can substantially reduce the amount of real-world learning (and mistakes) that robots make. A popular simulator for social navigation is the 2D CrowdNav simulator [6], designed for waypoint navigation in open spaces with the human policy defined by ORCA [7]. While CrowdNav has facilitated substantial DRL model development [9, 8, 10, 11], the cooperative nature of the ORCA policy creates a large sim-to-real gap. When the robot is visible to humans, the learned DRL policy is dangerously aggressive: the robot pushes humans out of its way to reach its goal [8]. Consequently, the invisible testbed—when the robot is invisible to the humans—was adopted as the benchmark standard. Although the invisible testbed reduces the DRL agent’s aggression, the problem becomes equivalent to decoupled planning where the DRL agent does not learn how its action will influence humans [8].
Beyond CrowdNav, other simulators use human motion models such as variations of the Social Forces Model (SFM) or behaviour graphs [12, 13]. Unfortunately, even the latest simulators struggle with human realism and often exhibit unsmooth human motions and enter deadlock scenarios. These deficiencies create a significant sim-to-real gap, where models trained in simulation often perform differently–—and usually worse–—when applied in the real world [14]. While simulators are essential tools for advancing DRL model architectures, the most accurate data for DRL social navigation is real-world data. By reducing the amount of training data required, we enable the direct training of DR-MPC in the real world, avoiding unnecessary inductive biases introduced by simulators.
@@
RL Social Navigation Models. Crowd navigation, a major sub-field of DRL social navigation, focuses on enabling a robot to navigate among humans in open areas. Significant progress has been made in incorporating machine learning advancements to enhance model reasoning about crowds in spatial and temporal dimensions.
@@
Early models analyzed the scene by considering humans individually [15, 16]. Then, approaches incorporated attention mechanisms to reason about the crowd as a whole, using learned attention scores to generate a crowd embedding for decision making [6, 17, 18]. Subsequent advances analyzed both human-robot and human-human interactions through graph neural networks (GNNs) [9, 18, 10, 19]. Most recently, models include temporal reasoning using spatio-temporal graphs, multihead attention mechanisms, and transformers [8, 11], enabling reasoning on the trajectory level.
@@
We do not focus on developing a novel architecture for waypoint-based crowd navigation; we incorporate state-of-the-art (SOTA) architectures for processing humans into DR-MPC.
@@
Residual DRL. Residual DRL integrates a user-supplied and a learned policy [20]. DRL learns corrective actions on top of the base controller. Learning speedup occurs because on initialization the model has an expected action equal to the base controller, which, assuming a suitable base controller, results in performance better than a random policy [20].
@@
To the best of our knowledge, residual DRL has not been applied to social navigation, but some works have attempted to incorporate classical control. Kästner et al. [21] learn a DRL policy that switches between different controllers, each optimized for an individual task. Another work, Semnani et al. [22] switch from a DRL policy to a force-based model when close to humans. While these approaches reduce DRL’s learning burden, the overall performance is limited by the base controllers. When the policy switches to a base controller, DRL can not optimize the individual controllers.
@@
In contrast, DR-MPC fully exploits the capabilities of MPC path tracking, which is an optimal behaviour when no humans are present. Like residual DRL, DR-MPC can replicate the MPC action or generate an action far from it. Unlike residual DRL, DR-MPC’s initial behaviour almost exactly follows MPC path tracking, the best performance without any prior human information. Additionally, DR-MPC learns to dynamically integrate or disregard the MPC action, leading to significantly faster training compared to residual DRL.
@@
Path Tracking. DRL is beneficial for path tracking when modeling difficulties arise, but in indoor environments with minimal disturbances and a robot that can be accurately modeled with unicycle kinematics, we can leverage MPC for path tracking using the formulation presented by Sehn et al. [23]. As a result, DRL does not need to explicitly learn path tracking, thereby reducing the overall learning complexity.
@@ III Decision process formulation
State Space. We construct our state $\mathcal{S}=\{\mathcal{S}_{\text{PT}},\mathcal{S}_{\text{HA}}\}$, where $\mathcal{S}_{\text{PT}}$ is the state information for path tracking and $\mathcal{S}_{\text{HA}}$ is the state information for human avoidance.
As in [8], we exclude human velocities due to the difficulty of estimating these quantities in the real world and assume a constant human radius. If at time $t$ there are $n_{t}$ visible humans: $\mathcal{S}_{\text{HA}}=\left\{\mathbf{v}^{t-H:t-1},\mathbf{r}^{t-H:t},\mathbf{q}^{t-H_{1}:t}_{1},\dots,\mathbf{q}^{t-H_{n_{t}}:t}_{n_{t}}\right\}$, where $\mathbf{v}^{t-H:t-1}$ is the robot’s past velocities from time $t-H$ to $t-1$, $\mathbf{r}^{t-H:t}$ is the robot’s past positions in the current robot frame from time $t-H$ to $t$, and $\mathbf{q}^{t-H_{i}:t}_{i}$ is the $i^{\text{th}}$ visible human’s past positions in the current robot frame from time $t-H_{i}$ to $t$, capped at length $H$.
We select a local path representation as in [24]. $\mathcal{S}_{\text{PT}}$ includes the path node closest to the robot, along with the preceding $L$ nodes and the following $F$ nodes. All nodes are transformed into the robot’s current reference frame.
Actions. The action space is a linear and angular velocity: $\mathbf{a}=\begin{bmatrix}v\hskip 5.69046pt\omega\end{bmatrix}$.
Rewards. Our reward function considers path advancement, path deviation, goal reaching, corridor collisions, small speeds, human collisions, and human disturbance:
@@
where rewards marked with an asterisk are terminal rewards. Path advancement: $r_{\text{pa}}=5\Delta s$ where $\Delta s$ is the arclength progress along the path. Deviation: $r_{\text{dev}}=0.5d_{\text{xy}}+0.03|d_{\theta}|$ where $d_{\text{xy}}$ is the Euclidean distance and $d_{\theta}$ is the angular offset from the closest point on the path. Goal: $r_{\text{goal}}^{*}=-5$ if the heading difference exceeds a threshold and 0 otherwise. Corridor collision: $r_{\text{cor-col}}^{*}=-10$ for colliding with the corridor. Minimal actuation: $r_{\text{act}}^{*}=-20$ if the sum of the robot’s past $H$ speeds falls below a threshold.
For human avoidance, we align our rewards with the two most important principles of human-robot interaction (HRI): safety and comfort [25]. For safety, human collision: $r_{\text{hum-col}}^{*}=-15$. For comfort, disturbance penalty [19]: $r_{\text{dist}}=-\sum_{i=1}^{n_{t}}\left(5.6|\Delta v_{\text{h}}^{i}|+3.5|\Delta\theta_{\text{h}}^{i}|\right)$, which penalizes the robot for causing changes in a human’s velocity ($\Delta v_{\text{h}}^{i}$) and heading direction ($\Delta\theta_{\text{h}}^{i}$). These changes are computed relative to the human’s velocity and heading at the previous time step.
@@ IV DR-MPC Policy Architecture
DR-MPC (Figure 3) consists of two main components: individual processing of $\mathcal{S}_{\text{PT}}$ and $\mathcal{S}_{\text{HA}}$, and the information fusion to generate a single action. We generate a latent embedding of the path using a Multi-layer Perceptron (MLP), and the MPC optimization from [23] generates $\mathbf{a}_{\text{MPC}}$.
To process $\mathcal{S}_{\text{HA}}$, we modify [10] to handle varying-length human trajectories for off-policy learning. We refer to this adapted architecture as the Human Avoidance Network (HAN); further details are provided in the appendix. The output of HAN is a ‘crowd embedding’ that is used to generate six candidate human avoidance actions ($\mathbf{a}_{\text{HA}}$), where each $\mathbf{a}^{\text{i}}_{\text{HA}}$ is positioned in a different cell of the action space. The human-avoidance action space is partitioned into six cells defined by two linear velocity bins $[v_{\text{min}},v_{\text{middle}}]$, $[v_{\text{middle}},v_{\text{max}}]$, and three angular velocity bins $[w_{\text{min}},w_{\text{lower}}]$, $[\omega_{\text{lower}},\omega_{\text{upper}}]$, $[\omega_{\text{upper}},\omega_{\text{max}}]$. We include multiple actions because often several viable actions exist for human avoidance. We found empirically that this design reduces learning time compared to having the model learn the diversity. We also found that using six actions strikes a good balance between sector granularity and the data required to adequately sample and explore each sector. The output of the MLP following HAN is the mean action within each cell. Using a predetermined variance that decays over time, we sample a Gaussian to get $\mathbf{a}^{\text{i}}_{\text{HA}}$; this standard formulation comes from [26].
@@
Unlike residual DRL, DR-MPC begins with near-MPC behaviour by biasing actions toward MPC, initializing $\beta=-0.8$ to suppress $\boldsymbol{\alpha}$. Note that initializing $\beta$ too low yields little qualitative difference and requires a lot of model updates to raise $\beta$ towards 0. We observe the model naturally learns to adjust $\beta$ towards 0 to be able to take non-MPC actions.
We train our model with Discrete Soft Actor-Critic (DSAC) with double average clipped Q-learning with Q-clip to properly backpropagate through $\mathbf{p}$ [28]. We augment DSAC with the entropy formulation from SAC to optimize the log-standard deviation that generates $\alpha^{i}_{\text{raw}}$ [27]. Figure 3 depicts the actor. The critic shares the same architecture up to and including the first MLP in the fusion stage; however, this MLP outputs the Q-value, and the action is included as input.